Circular No. 6394

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) J. Crovisier, E. Gérard, P. Colom, N. Biver, D. Bockelée-Morvan, and

H. Rauer, Observatoire de Paris-Meudon, report the detection of the 18-cm lines of OH with the Nançay radio telescope from the average of spectra observed on Apr. 12–19 and May 1–8: “The line areas at 1667 MHz are 0.028 ± 0.007 and 0.027 ± 0.007 Jy km s^{−1}, respectively. This would correspond to a water-production rate of 4.2 × 10^{28} molecules/s, using our usual OH excitation and space distribution model, which assumes outgassing from the nucleus. This rate is three times larger than that estimated for Apr. 7 from ultraviolet observations (IAUC 6376). This discrepancy may be due to shortcomings of the distribution model and might be explained by OH coming from icy grains.”

COMET C/1996 B2 (HYAKUTAKE) J. P. Pye and R. G. West, Leicester University; M. Harden and M.

Ricketts, Rutherford Appleton Laboratory; K. Dennerl, J. Englhauser, J. Schmitt, and J. Trümper, Max-Planck-Institut für Extraterrestrische Physik, Garching; and C. Lisse, M. Mumma, and R. Petre, NASA Goddard Space Flight Center, report the discovery of extreme-ultraviolet (EUV) emission from comet C/1996 B2 with the ROSAT Wide Field Camera (WFC), simultaneously with the x-ray measurements reported by Lisse et al. on IAUC 6373: “The WFC measurements were made with the S1A filter, nominal bandpass 90–206 eV (14–6 nm). An image can be found

at http://ledas-www.star.le.ac.uk/rosat-goc/comet/. Bright, dif-

fuse emission is seen, roughly coincident with the x-rays and sunward of the nucleus, and strongly time-variable in a similar manner. The peak EUV surface brightness is of order 0.0004 count s^{−1} arcmin^{−2} as measured, or of order 0.003 count s^{−1} arcmin^{−2} when corrected to WFC ‘at launch’ efficiency. The WFC images also show fainter diffuse emission extending beyond the edge of the HRI field-of-view (i.e., at ∼ 15′–40′ from the center of the comet nucleus), forming an arc around the bright central region. The ratio of WFC to HRI countrates shows that the spectrum in the EUV/xray band is rather ‘soft’ and steeply increasing towards low photon energies. This would appear to rule out soft-x-ray line-fluorescence as the dominant emission mechanism. For assumed spectral shapes of a power law or thermal bremsstrahlung emission, the ratio indicates a photon index in the approximate range 2.0–3.0, or a temperature of ∼ 0.1–0.4 keV.”

1996 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green